Find the value of y for the given value of x by first substituting the value of x into the equation and solving for y.

xy+x=-3; x=3


Answer by rfer(16322)

xy+x=-3
3y+3=-3
subtract 3 from both sides
3y+3-3=-3-3
3y=-6
divide both sides by 3
3y/3=-6/3
y=-2